# Wasserstein–Aitchison GAN for angular measures of multivariate extremes

### v2

@@ 2.3 Wasserstein Generative Adversial Networks
GANs have been introduced by Ian Goodfellow and co-authors in [goodfellow2014]. Initially, they have been proposed for synthetic image generation but since then, they have been applied to any kind of data, in particular tabular data which are of particular interest to statisticians [annoye2024]. See [bishop2024deep, Chapter 17] for a pedagogical introduction to GANs.
@@
A Wasserstein GAN (WGAN) architecture [arjovsky2017] follows this path to estimate $\mathsf{P}$ by considering $g_{\theta}$ to be a neural network and by measuring the difference in distribution between the real data distribution $\mathsf{P}$ and the law of $g_{\theta}(Z)$ via the Earth Mover’s Distance (EMD), which is a particular instance of the Wasserstein distance in optimal transport [villani2003].
@@
that is, for every $x\in\mathcal{X}$ except for a subset of Lebesgue measure zero. The equivalence between (13) and (14) is known as the Kantorovich–Rubinstein duality theorem, the proof of which is to be found in, e.g., [villani2003, Section 1.2].
@@
In practice there is no need for the neural network $D_{w}$ to be Lipschitz continuous as required in definition (14) of the EMD. Initially, in [arjovsky2017], the authors proposed to clip the weights, that is, forcing them to lie in a fixed compact space, typically $c,c]^{|\mathcal{S}_{D}|}$ for some small $c>0$. One may show that if the discriminator consists of a multilayer perceptron, which will always be the case here, its Lipschitz norm will be bounded by a finite constant depending only on $c$. Even though this constant may not be equal to one, replacing $\mathrm{Lip}_{1}$ by $\mathrm{Lip}_{C}$ for $C>0$ in (14) only amounts to multiplying the value of the resulting distance by $C>0$ and it does not play a role in the optimization of the generator.
@@
where $\mu$ is a probability measure on $\mathcal{X}$. The standard approach in [gulrajani2017] is to take $\mu$ as the law of the mixture $U\cdot X_{r}+(1-U)\cdot G_{\theta}(Z)$ where $X_{r}\sim\mathsf{P}$ follows the data distribution, $Z\sim\mathsf{P}_{Z}$ is random noise and $U$ is uniformly distributed on $[0,1]$ and independent of the rest. This choice is motivated by Proposition 1 in [gulrajani2017] and we follow this approach here.
@@ 2.4 Aitchison coordinates
A core idea of compositional data analysis [aitchison1982, Egozcue2003], mainly developed by John Aitchison (1926–2016), consists of transforming raw simplex data in $\mathbb{R}^{d}$ to coordinates with respect to an orthonormal basis of a certain (d1)(d-1)-dimensional subspace of $\mathbb{R}^{d}$. This makes interpretation more complicated, but having data on a full linear (d1)(d-1)-dimensional sample space instead of on the unit simplex in $\mathbb{R}^{d}$ can improve statistical and neural network modeling substantially. Since we are not all that concerned with interpretation in the WGAN architecture described in Section 2.3 but instead with obtaining the highest possible efficiency in producing accurate new multivariate extreme samples, we will apply this idea in our context.
Infinitely many orthogonal bases exist for the open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ equipped with its inner product space structure. Here, we work with the one proposed in Proposition 2 of [Egozcue2003]. Details of the underlying construction and more information on the Aitchison simplex, i.e., the unit simplex equipped with a certain vector space structure and an inner product, can be found in Appendix B. The essence of the matter is to map ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ to $\mathbb{H}=\{\bm{x}\in\mathbb{R}^{d}:\sum_{j=1}^{d}x_{j}=0\}$ by means of the centered logratio function ($\operatorname{clr}$) in (28). The set $\mathbb{H}$ is a (d1)(d-1)-dimensional linear space equipped with the standard Euclidean inner product, the structure of which carries over to ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ through the function $\operatorname{clr}$.
@@ 3 Combining GAN, Aitchison coordinates, and Extremes
The WA-GAN procedure in Algorithm 1 only uses the observations for which $R=|\bm{V}|_{1}$ is larger than $t=n/k_{1}$ so that, by (4), the number of observations used by the algorithm is asymptotically binomial with parameters $n$ and $(dk_{1})/n$. Here $k_{1}$ is a value to be chosen by the user, with larger $k_{1}$ meaning that more observations are used and hence the variance is smaller, and with smaller $k_{1}$ (hopefully) making the approximation in Assumption 1 more accurate. Finite sample bounds on the quality of this approximation are given in [clemencon2023concentration]. The relatively low effective sample size makes it important that the architectures of the generator and the discriminator are simple enough so that their parameters can be learned from a moderate number of examples. Those considerations are further explored in the numerical sections.
@@
   $\bm{W}_{i}^{*}\leftarrow\langle\operatorname{clr}(\bm{W}_{i}),\bm{e}_{i}\rangle$
@@
$V\leftarrow(\bm{e}_{1}\,|\,\ldots\,|\,\bm{e}_{d-1})\in\mathbb{R}^{d\times(d-1)}$
@@
  $L_{G}\leftarrow-\frac{1}{m}\sum_{i=1}^{m}D_{w}\left(G_{\theta}(\bm{z}_{i})\right)+\rho\left|\frac{1}{m}\sum_{i=1}^{m}\operatorname{softmax}\left(VG_{\theta}(\bm{z}_{i})\right)-\frac{\bm{1}_{d}}{d}\right|_{2}$
@@
Algorithm 1 is essentially the standard WGAN algorithm with gradient penalty of [gulrajani2017] applied to the coordinates of the large angles in an orthonormal basis of the Aitchison simplex. The ADAM optimizer which is used to update the weights of the generator and the discriminator at each step is a popular gradient descent algorithm with adaptive learning rate introduced in [kingma2015]. The function Adam() in the algorithm is one such gradient step and is the default optimizer proposed for the WGAN architecture with gradient penalty. Compared to the standard algorithm, one additional regularization term
$\rho\left|\frac{1}{m}\sum_{i=1}^{m}\operatorname{softmax}\left(VG_{\theta}(\bm{z}_{i})\right)-\frac{\bm{1}_{d}}{d}\right|_{2}$ (19)
is added to enforce the marginal constraint (3) in the generator in a soft way and thus to enforce sampling from a genuine angular measure.
Algorithm 1 focuses solely on the extremal dependence of $\bm{X}$ and is independent of any marginal assumptions besides continuity of the marginal distribution functions. Therefore, if one is interested in computing a quantity which only depends on $\Phi$ in Assumption 1, such as the coefficients $\theta_{J}$ introduced in the beginning of Section 4, this can be done independently of Algorithm 2. This is not the case for methods such as [Li-et-al] which directly evaluate sample quality on the scale of the MGP distribution. Note also that even though Algorithm 1 only considers the angular measure $\Phi$ with respect to the $L_{1}$-norm, a simple post-processing step allows for any norm on $\mathbb{R}^{d}$, as explained in Appendix C.
@@
The procedure in Algorithm 1 is based on data standardized to unit-Pareto margins, in accordance with Assumption 1. As in (18), we work with an empirical standardization $\widehat{F}_{j}$ for each margin $j=1,\ldots,d$, resulting in a procedure using only the marginal ranks of the original data and solely focusing on dependence, not requiring any assumption on the tail of $F_{j}$ as in Assumption 3. The same approach was followed in [einmahl2001nonparametric, einmahl2009maximum] and, more recently, in [clemencon2023concentration], for example. An alternative would be to use the empirical distribution function $\widehat{F}_{j}$ only up to a threshold $u_{j}<u_{*j}$ and fit a univariate GP distribution above this threshold, relying on Assumption 3. Doing so could improve the quality of the transformation to unit-Pareto margins, but at the price of making Algorithm 1 also require marginal assumptions. Furthermore, this route has already been explored in [einmahl1997estimating] and the same authors advocate in [einmahl2001nonparametric] for the pure nonparametric approach. Therefore, we stick to the rank-based procedure.
Some care needs is needed when sampling from $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$ for $t>1$ large, where $\bm{u}(t)$ is the vector of thresholds such that for each component separately, the excess probability is $1/t$. For example, a common situation is the case where $\bm{X}$ represents some positive multivariate risks, that is, $\bm{X}$ takes values in $0,\infty)^{d}$. If we directly make use of the formula $\bm{u}(t)+\bm{H}$, with estimated quantities, as suggested by formula (10), we risk generating “extreme” points for which some coordinates are negative. Our approach avoids this issue.
@@
$V\leftarrow(\bm{e}_{1}\,|\,\ldots\,|\,\bm{e}_{d-1})\in\mathbb{R}^{d\times(d-1)}$
@@
  $\bm{W}\leftarrow\operatorname{clr}^{-1}\left(VG_{\theta}(\bm{z})\right)$
@@ 4.1 Performance and computation
A first set of measures that we propose is based on the extremal coefficients  [beirlant2004, Section 8.2.7]. For any non-empty subset $J\subseteq\{1,\ldots,d\}$ with at least two elements $|J|\geq 2$, we define its extremal coefficient $\theta_{J}$ as
@@
The coefficients are estimated for the test set using the empirical angular measure in (21) and for the generative methods using the sampled realizations of $\Phi$. For the test set, the radial threshold $n/k$ is the same as for the training set. Typically, we will consider the bivariate coefficients, with $k=2$, as in [Boulagiem-et-al], and the trivariate coefficients, with $k=3$. This measure focuses on the extremal dependence structure only and aims to measure the quality of Algorithm 1.
A second measure that we propose, which also focuses on the tail of $\bm{X}$, aims at evaluating the output of Algorithm 2. Let $\bm{X}^{\mathcal{T}}$ denote a set of observations distributed like $\bm{X}$, which typically have not been used for training, and let $\bm{X}^{\mathcal{G}}_{\bm{u}}$ denote a set of observations generated by Algorithm 2 based on the threshold vector $\bm{u}=\bm{b}(n/k_{n})\in\mathbb{R}^{d}$. Then if the MGP approximation is accurate enough, we expect $\bm{X}_{\bm{u}}^{\mathcal{T}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bm{X}^{\mathcal{T}}\,|\,\bm{X}^{\mathcal{T}}\nleq\bm{u}$ to be close in distribution to $\bm{X}^{\mathcal{G}}_{\bm{u}}$. Hence, we propose to compute their difference in distribution using a simple extension of the EMD in Section 2.3 which is known as the $2$-Wasserstein distance [villani2003, Chapter 7]. If $n_{\mathcal{G}}$ and $n_{\mathcal{T}}$ denote, respectively, the size of the test set and the generated set of observations, the measure can be expressed as
@@
We can view $\pi_{ij}$ as the amount of mass transferred from $(\bm{X}_{\bm{u}}^{\mathcal{G}})_{i}$ to $(\bm{X}_{\bm{u}}^{\mathcal{T}})_{j}$. In a similar way as in definition (13) of the EMD, $W_{2}(\bm{X}_{\bm{u}}^{\mathcal{G}},\bm{X}^{\mathcal{T}}_{\bm{u}})$ measures, among all the possible ways to transport the empirical distribution of the vectors in $\bm{X}_{\bm{u}}^{\mathcal{G}}$ to the one of the vectors in $\bm{X}^{\mathcal{T}}_{\bm{u}}$, the cheapest way to do so if the cost of unitary transportation is equal to the squared Euclidean distance between points. In the case where we would have normally distributed quantities, this quantity would reduce to computing the well-known Fréchet inception distance, which is popularly used to assess the quality of generated images [heusel2017fid] using generative systems like the GAN, also used in [Li-et-al] to assess the quality of the MGP output of their algorithm. As we compare quantities that are far from being normally distributed, we think it makes more sense to compute the $2$-Wasserstein distance instead.
@@
For both the generator and the discriminator in Algorithm 1, we consider fully connected multilayer perceptrons with leaky rectified linear unit activation function [maas2013rectifier] which applies the map
@@
to each component of the linear transformation in a given layer, where $\alpha\in[0,1]$ is typically set to $0.01$. Taking $\alpha=0$ leads to the rectified linear unit activation function and $\alpha=1$ leads to a simple linear layer. We refer to [maas2013rectifier] for a comparison with other standard activation functions. Here, we consider $\alpha=0.01$ for any hidden layer while we take $\alpha=1$ for the final layer in each network. More recently, $\alpha$ has been considered as a learnable parameter in the optimization process [he2015delving] but this option is not considered in this work. Other architectures are possible if the data have some structure. For example, the authors of [Boulagiem-et-al] consider convolutions in their network as they work with spatial data on a grid, and this could be adapted to our framework too, but here we only consider basic multivariate data and keep this extension for future work.
@@
Inside the space of possible values for those hyperparameters, we perform a random search. Typically, we consider around $2000$ different models in this space for a given dataset. We train them with $n_{e}=5000$ epochs, since training them even longer did not significantly improve the performance, but greatly affects the computation time. Evaluation of the associated models is based on the extremal coefficient metric $E_{\overline{\theta}}(k)$ in (22) with $k=2$ and $k=3$. Validation and final performance assessment of the models are always performed on separate datasets.
@@
As reviewed in Section 2.1, numerous methods already exist in the rapidly growing field that combines generative approaches with extreme value theory, each with its own advantages and limitations. Therefore, an exhaustive comparison with all existing methods is not feasible, even less so as the software underlying the numerical results is often not available from the papers. Instead, we choose to compare our approach with versions of two recent methods: HTGAN from [Girard-et-al] and GPGAN from [Li-et-al]. We discuss our specific implementations in Appendix D in the supplement.
@@
The code used to produce the results of this section was written in Python—in particular, the PyTorch [PyTorch2019] framework was used to train the generative models—and in R [R2023], which was mainly used for data preprocessing and computation of the performance measures associated with the generated data, notably via the transport [Rtransport2024] package for the computation of the Wasserstein distance $W_{2}$ in (23).

### v3

@@ 2.2 Wasserstein Generative Adversial Networks
A Wasserstein GAN (WGAN) architecture [5] follows this path to estimate $\mathsf{P}$ by considering $g_{\theta}$ to be a neural network and by measuring the difference in distribution between the real data distribution $\mathsf{P}$ and the law of $g_{\theta}(Z)$ via the Earth Mover’s Distance (EMD), which is a particular instance of the Wasserstein distance in optimal transport [56].
@@
that is, for every $x\in\mathcal{X}$ except for a subset of Lebesgue measure zero. The equivalence between (13) and (14) is known as the Kantorovich–Rubinstein duality theorem, the proof of which is to be found in, e.g., [56, Section 1.2].
@@
In practice there is no need for the neural network $D_{w}$ to be Lipschitz continuous as required in definition (14) of the EMD. Initially, in [5], the authors proposed to clip the weights, that is, forcing them to lie in a fixed compact space, typically $c,c]^{|\mathcal{S}_{D}|}$ for some small $c>0$. One may show that if the discriminator consists of a multilayer perceptron, which will always be the case here, its Lipschitz norm will be bounded by a finite constant depending only on $c$. Even though this constant may not be equal to one, replacing $\mathrm{Lip}_{1}$ by $\mathrm{Lip}_{C}$ for $C>0$ in (14) only amounts to multiplying the value of the resulting distance by $C>0$ and it does not play a role in the optimization of the generator.
@@
where $\mu$ is a probability measure on $\mathcal{X}$. The standard approach in [27] is to take $\mu$ as the law of the mixture $U\cdot X_{r}+(1-U)\cdot G_{\theta}(Z)$ where $X_{r}\sim\mathsf{P}$ follows the data distribution, $Z\sim\mathsf{P}_{Z}$ is random noise and $U$ is uniformly distributed on $[0,1]$ and independent of the rest. This choice is motivated by Proposition 1 in [27] and we follow this approach here.
@@ 2.3 Aitchison coordinates
A core idea of compositional data analysis [1, 18], mainly developed by John Aitchison (1926–2016), consists of transforming raw simplex data in $\mathbb{R}^{d}$ to coordinates with respect to an orthonormal basis of a certain (d1)(d-1)-dimensional subspace of $\mathbb{R}^{d}$. This makes interpretation more complicated, but having data on a full linear (d1)(d-1)-dimensional sample space instead of on the unit simplex in $\mathbb{R}^{d}$ can improve statistical and neural network modeling substantially. Since we are not all that concerned with interpretation in the WGAN architecture described in Section 2.2 but instead with obtaining the highest possible efficiency in producing accurate new multivariate extreme samples, we will apply this idea in our context.
Infinitely many orthogonal bases exist for the open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ equipped with its inner product space structure. Here, we work with the one proposed in Proposition 2 of [18]. Details of the underlying construction and more information on the Aitchison simplex, i.e., the unit simplex equipped with a certain vector space structure and an inner product, can be found in Appendix B. The essence of the matter is to map ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ to $\mathbb{H}=\{\bm{x}\in\mathbb{R}^{d}:\sum_{j=1}^{d}x_{j}=0\}$ by means of the centered logratio function ($\operatorname{clr}$) in (28). The set $\mathbb{H}$ is a (d1)(d-1)-dimensional linear space equipped with the standard Euclidean inner product, the structure of which carries over to ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ through the function $\operatorname{clr}$. Recall that for a vector $\bm{x}\in\mathbb{R}^{d}$, $\operatorname{softmax}(\bm{x})$ is the element of ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ defined by
$\operatorname{softmax}(\bm{x})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\frac{\exp(x_{j})}{\sum_{i=1}^{d}\exp(x_{i})}\right)_{j=1}^{d}.$
@@ 3 Combining GAN, Aitchison coordinates, and Extremes
The WA-GAN procedure in Algorithm 1 only uses the observations for which $R=|\bm{V}|_{1}$ is larger than $t=n/k_{1}$ so that, by (4), the number of observations used by the algorithm is asymptotically binomial with parameters $n$ and $(dk_{1})/n$. Here $k_{1}$ is a value to be chosen by the user, with larger $k_{1}$ meaning that more observations are used and hence the variance is smaller, and with smaller $k_{1}$ (hopefully) making the approximation in Assumption 1 more accurate. Finite sample bounds on the quality of this approximation are given in [12]. The relatively low effective sample size makes it important that the architectures of the generator and the discriminator are simple enough so that their parameters can be learned from a moderate number of examples. Those considerations are further explored in the numerical sections.
@@
   $\bm{W}_{i}^{*}\leftarrow(\langle\operatorname{clr}(\bm{W}_{i}),\bm{e}_{j}\rangle)_{j=1}^{d-1}$
@@
$M\leftarrow(\bm{e}_{1}\,|\,\ldots\,|\,\bm{e}_{d-1})\in\mathbb{R}^{d\times(d-1)}$
@@
  $L_{G}\leftarrow-\frac{1}{m}\sum_{i=1}^{m}D_{w}\left(G_{\theta}(\bm{z}_{i})\right)+\rho\left|\frac{1}{m}\sum_{i=1}^{m}\operatorname{softmax}\left(MG_{\theta}(\bm{z}_{i})\right)-\frac{\bm{1}_{d}}{d}\right|_{2}^{2}$
@@
Algorithm 1 is essentially the standard WGAN algorithm with gradient penalty of [27] applied to the coordinates of the large angles in an orthonormal basis of the Aitchison simplex. The ADAM optimizer which is used to update the weights of the generator and the discriminator at each step is a popular gradient descent algorithm with adaptive learning rate introduced in [35]. The function Adam() in the algorithm is one such gradient step and is the default optimizer proposed for the WGAN architecture with gradient penalty. Compared to the standard algorithm, one additional regularization term
$\rho\left|\frac{1}{m}\sum_{i=1}^{m}\operatorname{softmax}\left(MG_{\theta}(\bm{z}_{i})\right)-\frac{\bm{1}_{d}}{d}\right|_{2}^{2},$ (19)
where $M$ is the matrix defined in Algorithm 1, is added to enforce the marginal constraint (3) in the generator in a soft way and thus to enforce sampling from a genuine angular measure. The expression in (19) penalizes the (empirical) deviations from the marginal conditions in (3) by forcing the squared Euclidean norm of their differences to be close to zero. Note that $MG_{\theta}(\bm{z}_{i})$ is the point in $\mathbb{H}$ with coordinates $G_{\theta}(\bm{z}_{i})$ with respect to the orthonormal basis $\{\bm{e}_{1},\ldots,\bm{e}_{d-1}\}$; the $\operatorname{softmax}$ function maps this point to the unit simplex. The regularization term then encourages the empirical distribution of the $m$ points on the unit simplex obtained in this way to the moment constraints in (3).
Algorithm 1 focuses solely on the extremal dependence of $\bm{X}$ and is independent of any marginal assumptions besides continuity of the marginal distribution functions. Therefore, if one is interested in computing a quantity which only depends on $\Phi$ in Assumption 1, such as the coefficients $\theta_{J}$ introduced in the beginning of Section 4, this can be done independently of Algorithm 2. This is not the case for methods such as [37] which directly evaluate sample quality on the scale of the MGP distribution. Note also that even though Algorithm 1 only considers the angular measure $\Phi$ with respect to the $L_{1}$-norm, a simple post-processing step allows for any norm on $\mathbb{R}^{d}$, as explained in Appendix C.
@@
The procedure in Algorithm 1 is based on data standardized to unit-Pareto margins, in accordance with Assumption 1. As in (18), we work with an empirical standardization $\widehat{F}_{j}$ for each margin $j=1,\ldots,d$, resulting in a procedure using only the marginal ranks of the original data and solely focusing on dependence, not requiring any assumption on the tail of $F_{j}$ as in Assumption 3. The same approach was followed in [19, 21] and, more recently, in [12], for example. An alternative would be to use the empirical distribution function $\widehat{F}_{j}$ only up to a threshold $u_{j}<u_{*j}$ and fit a univariate GP distribution above this threshold, relying on Assumption 3. Doing so could improve the quality of the transformation to unit-Pareto margins, but at the price of making Algorithm 1 also require marginal assumptions. Furthermore, this route has already been explored in [20] and the same authors advocate in [19] for the pure nonparametric approach. Therefore, we stick to the rank-based procedure.
Some care is needed when sampling from $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$ for $t>1$ large, where $\bm{u}(t)$ is the vector of thresholds such that for each component separately, the excess probability is $1/t$. For example, a common situation is the case where $\bm{X}$ represents some positive multivariate risks, that is, $\bm{X}$ takes values in $0,\infty)^{d}$. If we directly make use of the formula $\bm{u}(t)+\bm{H}$, with estimated quantities, as suggested by formula (10), we risk generating “extreme” points for which some coordinates are negative. Our approach avoids this issue.
@@
$M\leftarrow(\bm{e}_{1}\,|\,\ldots\,|\,\bm{e}_{d-1})\in\mathbb{R}^{d\times(d-1)}$
@@
  $\bm{W}\leftarrow\operatorname{clr}^{-1}\left(MG_{\theta}(\bm{z})\right)$
@@ 4.1 Performance and computation
A first set of measures that we propose is based on the extremal coefficients  [7, Section 8.2.7]. For any non-empty subset $J\subseteq\{1,\ldots,d\}$ with at least two elements $|J|\geq 2$, we define its extremal coefficient $\theta_{J}$ as
@@
The coefficients are estimated for the test set using the empirical angular measure in (21) and for the generative methods using the sampled realizations of $\Phi$. For the test set, the radial threshold $n/k$ is the same as for the training set. Typically, we will consider the bivariate coefficients, with $k=2$, as in [11], and the trivariate coefficients, with $k=3$. This measure focuses on the extremal dependence structure only and aims to measure the quality of Algorithm 1.
A second measure that we propose, which also focuses on the tail of $\bm{X}$, aims at evaluating the output of Algorithm 2. Let $\bm{X}^{\mathcal{T}}$ denote a set of observations distributed like $\bm{X}$, which typically have not been used for training, and let $\bm{X}^{\mathcal{G}}_{\bm{u}}$ denote a set of observations generated by Algorithm 2 based on the threshold vector $\bm{u}=\bm{b}(n/k_{n})\in\mathbb{R}^{d}$. Then if the MGP approximation is accurate enough, we expect $\bm{X}_{\bm{u}}^{\mathcal{T}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bm{X}^{\mathcal{T}}\,|\,\bm{X}^{\mathcal{T}}\nleq\bm{u}$ to be close in distribution to $\bm{X}^{\mathcal{G}}_{\bm{u}}$. Hence, we propose to compute their difference in distribution using a simple extension of the EMD in Section 2.2 which is known as the $2$-Wasserstein distance [56, Chapter 7]. If $n_{\mathcal{G}}$ and $n_{\mathcal{T}}$ denote, respectively, the size of the test set and the generated set of observations, the measure can be expressed as
@@
We can view $\pi_{ij}$ as the amount of mass transferred from $(\bm{X}_{\bm{u}}^{\mathcal{G}})_{i}$ to $(\bm{X}_{\bm{u}}^{\mathcal{T}})_{j}$. In a similar way as in definition (13) of the EMD, $W_{2}(\bm{X}_{\bm{u}}^{\mathcal{G}},\bm{X}^{\mathcal{T}}_{\bm{u}})$ measures, among all the possible ways to transport the empirical distribution of the vectors in $\bm{X}_{\bm{u}}^{\mathcal{G}}$ to the one of the vectors in $\bm{X}^{\mathcal{T}}_{\bm{u}}$, the cheapest way to do so if the cost of unitary transportation is equal to the squared Euclidean distance between points. In the case where we would have normally distributed quantities, this quantity would reduce to computing the well-known Fréchet inception distance, which is popularly used to assess the quality of generated images [30] using generative systems like the GAN, also used in [37] to assess the quality of the MGP output of their algorithm. As we compare quantities that are far from being normally distributed, we think it makes more sense to compute the $2$-Wasserstein distance instead.
@@
For both the generator and the discriminator in Algorithm 1, we consider fully connected multilayer perceptrons with leaky rectified linear unit activation function [38] which applies the map
@@
to each component of the linear transformation in a given layer, where $\alpha\in[0,1]$ is typically set to $0.01$. Taking $\alpha=0$ leads to the rectified linear unit activation function and $\alpha=1$ leads to a simple linear layer. We refer to [38] for a comparison with other standard activation functions. Here, we consider $\alpha=0.01$ for any hidden layer while we take $\alpha=1$ for the final layer in each network. More recently, $\alpha$ has been considered as a learnable parameter in the optimization process [29] but this option is not considered in this work. Other architectures are possible if the data have some structure. For example, the authors of [11] consider convolutions in their network as they work with spatial data on a grid, and this could be adapted to our framework too, but here we only consider basic multivariate data and keep this extension for future work.
@@
Inside the space of possible values for those hyperparameters, we perform a random search. Typically, we consider around $2000$ different models in this space for a given dataset. We train them with $n_{e}=5000$ epochs, since training them even longer did not significantly improve the performance, but greatly affects the computation time. Evaluation of the associated models is based on the extremal coefficient metric $E_{\overline{\theta}}(k)$ in (22) with $k=2$ and $k=3$. The validated model is chosen to be the one with the lowest value of $\left(E_{\overline{\theta}}(2)+E_{\overline{\theta}}(3)\right)/2$ among the $2000$ models. This validation is performed on a separate dataset from the one used for training or testing.
@@
As reviewed in Section 1, numerous methods already exist in the rapidly growing field that combines generative approaches with extreme value theory, each with its own advantages and limitations. Therefore, an exhaustive comparison with all existing methods is not feasible, even less so as the software underlying the numerical results is often not available from the papers. Instead, we choose to compare our approach with versions of two recent methods: HTGAN from [25] and GPGAN from [37]. We discuss our specific implementations in Appendix D in the supplement.
@@
The code used to produce the results of this section was written in Python—in particular, the PyTorch [44] framework was used to train the generative models—and in R [46], which was mainly used for data preprocessing and computation of the performance measures associated with the generated data, notably via the transport [53] package for the computation of the Wasserstein distance $W_{2}$ in (23).